OpenStudy (anonymous):

A pyramid has a square base with side s. The height of the pyramid is 2/3 that of its side. What is the expression for the volume of the pyramid?

OpenStudy (anonymous):

V = 1/3 * s^2 * (2/3)s = (2/9)s^3